Suppose John have R1 800 to spend on two goods: shoes, and shirts and that the price of a pair of shoes is R120 and the price of shirt is R100. (i). Given total income and prices of two goods, how many pairs of shoes will John obtain if he decides to use all of his income on shoes and how many units of shirts will he afford if he decides to buy shirts only?What will happen to John’s budget line if the price of shoes (measured on horizontal axis rise from R120 to R180? Explain by stating numerical figures.


1
Expert's answer
2021-07-13T11:58:16-0400

Solution:

i.).I = PxX + PyY

1800 = 120X + 100Y

Income on pair of shoes only = "\frac{1800}{120}" = 15 pairs

John will obtain 15 pairs of shoes if he decides to use all his income on shoes.

 

Income on shirts only = "\frac{1800}{100}" = 18 units

John will obtain 18 units of shirts if he decides to use all his income on shirts.

 

John’s budget line will shift inward, pivoting from the shoe's vertical intercept. The horizontal x-intercept will shift in.

This is because John will not be able to afford more shoes due to an increase in price and the demand for shoes will decrease while maintaining the same demand for shirts.


The initial demand for shoes was = "\frac{1800}{120}" = 15 pairs


New demand after increase in price = "\frac{1800}{180}" = 10 pairs


Therefore, the demand for shoes will decrease by 5 pairs as a result of its price increase.